# When Do Credal Sets Stabilize? Fixed-Point Theorems for Credal Set Updates

### v1

@@ 3. Main Results
The reader may ask themselves why, in Corollary 2.2, we chose to prove a generalization of the Boyd-Wong fixed point theorem; in other words, whether we could have showed something more general.
First, let us discuss the conditions on $\mathcal{X}$, and recall that, by Remark 2, we only need it to be Polish. Notice that if $\mathcal{X}$ is not separable, then $\Delta_{\mathcal{X}}$ may not be metrizable, which in turn means that we cannot talk about the Hausdorff metric $d_{H}$ on $\mathscr{C}$, let alone about Hausdorff convergence. Also, if $\mathcal{X}$ is not complete, then there might be an orbit $(\mathcal{P}_{n})$ of some $\mathcal{P}_{0}\in\mathscr{C}$ that has an accumulation point, but does not converge to it.
@@
Example 3 (CBDL, continued).
Let us continue working in the setting of Example 2. For simplicity, we restrict to priors in $\Delta_{\Theta}^{++}$, i.e. those with densities bounded away from 0 and $\infty$ with respect to $\mu$.
Remark 4 (Why the restriction is minimal).
For any $\theta_{1}\neq\theta_{2}$, $T_{k}(\delta_{\theta_{1}})=\delta_{\theta_{1}}$ and $T_{k}(\delta_{\theta_{2}})=\delta_{\theta_{2}}$, so no metric that separates atoms can give a global contraction constant <1<1 on $\Delta_{\Theta}$. Excluding atoms, and therefore working with $\Delta_{\Theta}^{++}$, is therefore necessary.
Proposition 3.
Fix the measurable set $E\subset\mathcal{Y}$ representing the gathered evidence. Suppose that (I) and (II) still hold, but require the following strengthening of the condition in (III) for the likelihoods $L_{k}\in\text{ex}\mathcal{L}$,
  1. (III’)
    For all $L_{k}\in\text{ex}\mathcal{L}$, there exist constants $0<\alpha_{k}\leq\beta_{k}<\infty$ such that
    $\alpha_{k}\leq\ell_{k}(E\mid\theta)\leq\beta_{k},\quad\forall\theta\in\Theta.$
Then, we can find a function $\psi$ that satisfies (i)-(ii) in Corollary 2.2. It is
$t\mapsto\psi(t)\coloneqq\sup\left\{d_{H}(f(\mathcal{P}),f(\mathcal{Q})):d_{H}(\mathcal{P},\mathcal{Q})\leq t\right\}.$
Proposition 3 implies that, under conditions (I), (II), (III’), we can apply Corollary 2.2 to conclude that the CBDL updating technique for a prior FGCS $\mathcal{P}$ has a unique fixed point $\mathcal{P}_{\star}$, and that the sequence $(\mathcal{P}_{n})$ of the orbits of $\mathcal{P}$ under $f$ converges to $\mathcal{P}_{\star}$.
In addition, the proof of Proposition 3 (which can be found in Appendix A.4) gives us a nice intuition. If two prior FGCSs are close, their updates get even closer, capturing a kind of informational convergence under CBDL and assumptions (I), (II), (III’).
$\underline{g}_{k}(P^{\text{ex}})(A)\coloneqq\inf_{E\in\mathcal{E}}\frac{P^{\text{ex}}(A)L_{k}(E\mid A)}{L_{k}(E)},$ (3)
where $\mathcal{E}\subset 2^{\mathcal{Y}}$ is a collection of subsets of $\mathcal{Y}$ whose elements are measurable and such that the ratio is well defined. [^5] More precisely, $\mathcal{E}\subset\mathcal{F}_{\mathcal{Y}}$, the sigma-algebra on $\mathcal{Y}$. Notice that the evidence $E$ depends on $A\subset\Theta$, and we write $E\equiv E(A)$ for notational convenience.
Condition (3) tells us two things. First, instead of fixing $E$, we fix a class $\mathcal{E}$ of measurable sets that are associated with the experiment at hand. That is, no matter what step $n\in\mathbb{Z}_{+}$ we are considering, we know that the evidence $E$ that we may gather belongs to $\mathcal{E}$. To borrow terminology from Walley, (1991, Section 2.1.2), we say that the elements of $\mathcal{E}$ are the pragmatic possible outcomes of the experiment. For example, if we consider a coin tossing, we may posit that $\mathcal{E}=2^{\{H,T\}}$, where $H$ denotes heads and $T$ denotes tails, instead of considering the power set of all apparently possible outcomes (Caprio and Mukherjee,, 2023, Section 3). These include e.g. coin landing on its edge, coin breaking into pieces on landing, coin disappearing down a crack in the floor.
  1. (IV’)
[…]
It is then easy to see, then, that, under conditions (I), (II), (III’), (IV’), the $\underline{g}_{k}$’s are continuous and the $\underline{g}_{k}(P^{\text{ex}})$’s are countably additive probabilities, so all the results we derived previously continue to hold.
More in general, our results hold for modifications to the CBDL updating mechanism that ensure continuity of (a modification of) the $g_{k}$’s, and that (a modification of) the ${g}_{k}(P^{\text{ex}})$’s are Kolmogorovian probabilities.  $\triangle$
From Example 3, it is clear that sometimes studying the “orbit” of a credal set $\mathcal{P}_{0}$ under varying functions $f_{i}$, $i\in\{1,\ldots,n\}$, as $n\rightarrow\infty$, is of great interest. Let us address such a case.
Suppose that we collect evidence in the form of sets $E_{1},\ldots,E_{m}\in\mathcal{F}_{\mathcal{Y}}$, $m\in\mathbb{N}_{\geq 2}$, where $\mathcal{F}_{\mathcal{Y}}$ denotes the $\sigma$-algebra of the evidence space $\mathcal{Y}$, possibly different from $\mathcal{X}$. Call $\breve{E}\coloneqq\cup_{i=1}^{m}E_{i}$, and put $\mathcal{E}=2^{\breve{E}}$. Consider our updating functions $f_{E}:\mathscr{C}\rightarrow\mathscr{C}$, this time indexed by some $E\in\mathcal{E}$, and let $\overline{f}\coloneqq\sup_{E\in\mathcal{E}}f_{E}$ and $\underline{f}\coloneqq\inf_{E\in\mathcal{E}}f_{E}$. These are well defined because $\overline{f}=\overline{\text{CH}(\cup_{E\in\mathcal{E}}f_{E})}$, and we assume $\underline{f}=\cap_{E\in\mathcal{E}}f_{E}\neq\emptyset$, so that $\underline{f}(\mathcal{P})\in\mathscr{C}$, for all $\mathcal{P}$. It is easy to see, then, that $\underline{f}(\mathcal{P}_{0})\subseteq f_{E}(\mathcal{P}_{0})\subseteq\overline{f}(\mathcal{P}_{0})$, for all $E\in\mathcal{E}$ and all $\mathcal{P}_{0}\in\mathscr{C}$.
Now, fix any $n\in\mathbb{N}$, and define $\bigcirc_{i=1}^{n}f_{E_{i}}\coloneqq f_{E_{n}}\circ\cdots\circ f_{E_{1}}$, where $\{E_{i}\}_{i=1}^{n}\subset\mathcal{E}$. We have the following.
  • The conditions in Corollary 2.2 hold for both $\underline{f}$ and $\overline{f}$.
Two remarks are in order. First, as we can see from the statement of Proposition 4, we need to ensure that the Hausdorff limit $\lim_{n\rightarrow\infty}\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})$ exists on a case by case basis. In other words, we need more mathematical structure around how the functions $f_{E}$ update our initial credal set $\mathcal{P}_{0}$; We cannot gauge their limiting behavior (when composed with one another) at the level of generality that we work on in Proposition 4. Two simple sufficient conditions are (a) that the $f_{E_{i}}$’s are all strict Hausdorff contraction whose Lipschitz constants multiply to 0, or (b) that the images of the $f_{E_{i}}$’s form a nested chain whose diameters shrink to 0.
Second, alas CBDL does not satisfy the order-preserving condition, as (generalized) Bayes’ updating is well known to dilate (Seidenfeld and Wasserman,, 1993). Some mechanisms that satisfy the order-preserving condition can be found in Caprio and Seidenfeld, (2023).
Notice that the only information we get from Corollary 2.2 about the convergence rate of $(\mathcal{P}_{n})$ is that it is sublinear and governed by $\psi$. For example, if $\psi$ satisfies additional properties—e.g. Lipschitzianity or $\psi(t)=qt$, with $q\in(0,1)$—the rate becomes tractable, as the setting becomes similar to Banach’s. It is also worth mentioning that there is an active literature studying the convergence speed for Boyd-Wong-type contractions under extra assumptions on the structure of $\psi$ (Gupta and Mansotra,, 2023; Gautam and Kaur,, 2023; Singh et al.,, 2022; Das et al.,, 2024).
Example 4 (CBDL, continued).
Before concluding this section, let us find the rate of convergence for the sequence $(\mathcal{P}_{n})$ in Example 3. From the proof of Proposition 3 in Appendix A.4, the control function that we found is
$\psi(t)\coloneqq\sup\left\{d_{H}(f(\mathcal{P}),f(\mathcal{Q})):d_{H}(\mathcal{P},\mathcal{Q})\leq t\right\},$
and we denote by $(\mathcal{P}_{n})$ the sequence of orbits of a prior FGCS $\mathcal{P}$ under $f$. Let $d_{n}\coloneqq d_{H}(\mathcal{P}_{n},\mathcal{P}_{\star})$, where $\mathcal{P}_{\star}$ is the unique fixed point of $f$, as per Corollary 2.2, and $d_{H}$ denotes the Hausdorff metric.
Then by Proposition 6 in Appendix A.4, the control function $\psi$ satisfies the linear bound
Hence,
and therefore
$d_{n}\leq\tau^{n}d_{0}\qquad(n\geq 0).$
In particular, the orbit converges to $\mathcal{P}_{\star}$ at a geometric rate,
$d_{H}(\mathcal{P}_{n},\mathcal{P}_{\star})=\mathcal{O}(\tau^{n}).$
Assumptions used. This bound relies on the standing assumptions from Example 3: priors in $\Delta_{\Theta}^{++}$ and likelihoods satisfying (III’).  $\triangle$
4. Illustration with Finitely Generated Credal Sets
Figure 2. Illustration of finitely generated credal sets reaching their fixed points. Starting with a credal set generated by finitely many extreme distributions, we repeatedly apply the updating map $f$. With each iteration, the distance between the parameters of the distributions located at the extreme corners of the set decreases. This contraction effect continues until the successive distances vanish, at which point the credal set collapses to its fixed point.
We provide a simple illustration of our main result (Theorem 1) using the following set-up. We assume the following data-generating process:
$\theta_{\star}$ $=0$
$X_{1}\dots,X_{n}$ $\sim\operatorname{N}(\theta_{\star},1)$
We generate our prior finitely generated credal set $\mathcal{P}_{0}$ by setting
$P_{j}(\theta)=\operatorname{N}(\theta\mid\mu_{j},\tau^{2}_{j}),\quad j\in\{1,2,3\},$
and $\mathcal{P}_{0}=\operatorname{CH}(\{P_{j}\}_{j=1}^{3})$. Following the notation from Example 2, we can define the updating procedure of this credal set by $f$, where
$f(\mathcal{P}):=\operatorname{CH}(\varphi(\operatorname{ex}(\mathcal{P})),$
and $\varphi(\operatorname{ex}(\mathcal{P}))$ denotes the elementwise Bayes update with the given evidence $X_{1},\dots,X_{n}$ to each prior $P_{1},P_{2},P_{3}$. As a Gaussian prior on the mean is conjugate to a Gaussian likelihood with known variance, we have a closed form for the posterior finitely generated credal set, given by
$P_{j}(\theta\mid X_{1},\dots X_{n})=\operatorname{N}(\theta\mid\tilde{\mu}_{j},\tilde{\tau}_{j}^{2}),\quad j\in\{1,2,3\},$
where
$\tilde{\tau}_{j}^{2}=\left(\frac{1}{\tau_{j}^{2}}+n\right)^{-1},\quad\tilde{\mu}_{j}=\tilde{\tau_{j}}^{2}\left(\frac{\mu_{j}}{\tau_{j}^{2}}+\sum_{i=1}^{n}X_{i}\right).$
For the illustration, we set $\mu_{1},\mu_{2},\mu_{3}=(2,3,-2)$ and $\tau_{1}^{2},\tau_{2}^{2},\tau_{3}^{2}=(1,2,2)$. To establish that the credal set $\mathcal{P}$ reaches a fixed point under repeated application of the updating map $f$, it is sufficient to examine only the extreme distributions. Equivalently, instead of tracking full distributions, we only need to verify whether the parameters of the Gaussians $(\mu_{i},\tau_{i}^{2})$ converge to fixed values. If each sequence of parameters stabilises, then the whole credal set stabilises.
Figure 2 illustrates the result of repeatedly applying $f$ to the prior finitely generated credal set $\mathcal{P}$. We pack the $t^{\text{th}}$ round of parameters $(\theta_{j,t},\tau^{2}_{j,t})$ together and denote it as $\xi_{t}^{(j)}$. We plot the averaged consecutive parameter differences $\frac{1}{3}\sum_{j=1}^{3}\|\xi^{(j)}_{t}-\xi^{(j)}_{t-1}\|_{2}^{2}$ on the y-axis and the number of rounds in the x-axis. As illustrated, the difference quickly converges to 0, meaning the credal sets reach their fixed point.
5. Discussion
In this work, we made a first stride towards the study of fixed point results for credal sets. In particular, we derived credal-sets-versions of the celebrated Edelstein and Boyd-Wong fixed point theorems.
Our findings open up many interesting questions, which will be the topic of future research. First, although intuitive, it remains to be explicitly shown that (Bayesian) posterior consistency/concentration can be framed as special cases of our results. Correctly specifying the likelihood, and then applying Bayes’ rule iteratively, will result in reaching a fixed point being the Dirac distribution at the right parameter.
Second, we plan to derive an axiomatization of the desirable properties of an updating mechanism, a goal in line with the recent literature in imprecise opinion pooling (Elkin and Pettigrew,, 2025).
Third, notice that, although our examples were based on CBDL, we do not confine our attention to the Bayesian paradigm alone. This means that our results are useful to any credal set learning paradigm. Consider the case where we start with a set of predictors (initialising parameters randomly or having different hyper parameter, as it is done e.g. in deep ensemble) and then, after collecting evidence, we either update the models separately, or “trim” them. That is, only a few are kept, e.g. those which clear a certain threshold for being the “correct ones”. In the future, we plan to study the structure needed to place on these kinds of threshold-based update rules in addition to the standard learning procedure, to have fixed points, and to be able to reach them.
Finally, we plan to inspect how our theorems can be put to use to study the existence of fixed point theorems for lower probabilities in the context of IIPM (Chau et al., 2025a, ), a generalisation that take also into account epistemic uncertainty, of the well-studied field of Integral Probability Metrics.
Appendix A Additional theorems and Proofs
A.1. Proof for Theorem 1
@@
Brouwer-Schauder-Tychonoff fixed-point theorem (Aliprantis and Border,, 2006, Theorem 17.56) guarantees a point $h_{*}\in K$ with $\tilde{f}(h_{*})=h_{*}$, since $K$ is compact and convex in the locally convex topological (Hausdorff) vector space $Z$.
@@
Let $P_{*}\coloneqq J^{-1}(h_{*})$. Then $f(P_{*})=J^{-1}\bigl(\tilde{f}(J(P_{*}))\bigr)=P_{*}$, so $P_{*}$ is a fixed point of $f$. Finally, the collection of fixed points of $f$, $\mathrm{Fix}(f)=\{P\in\mathscr{C}:d_{H}(P,f(P))=0\}$, is closed in compact $\mathscr{C}$, hence compact. ∎
A.2. Proof for Proposition 2
@@
Notice now that, at the beginning of the proof, we picked a generic prior $P^{\text{ex}}\in\text{ex}\mathcal{P}$ and a generic likelihood $L_{k}\in\text{ex}\mathcal{L}$. This means that the functions $A\mapsto g_{k}(P^{\text{ex}})(A)$ are continuous, for all $k$ and all $P^{\text{ex}}$. Then, since $\Delta_{\Theta}$ is locally convex (as a result of $\Theta$ being compact metric), we conclude that $f$ is Hausdorff continuous by Aliprantis and Border, (2006, Theorem 17.37). ∎
A.3. Proof for Remark 2
@@
Thus, $C\in\mathscr{C}$, proving that $\mathscr{C}$ is a closed subset of $\mathscr{F}$. Since $\mathscr{F}$ is complete under $d_{H}$, and $\mathscr{C}$ is a closed subset of $\mathscr{F}$, it follows that $\mathscr{C}$ is complete under $d_{H}$. ∎
A.4. Proof for Proposition 3
@@
Multiplication by $\ell_{k}(E\mid\cdot)$ defines a positive linear map
$M_{k}:p\mapsto\ell_{k}(E\mid\cdot)p$
on the positive cone of $L^{1}(\mu)$; the normalization $p\mapsto p/\int p\text{d}\mu$ is projective (leaves $d_{\mathrm{H}}$ unchanged). Since $\alpha_{k}\leq\ell_{k}\leq\beta_{k}$ a.e., the projective diameter of $M_{k}$ equals $\mathscr{D}(M_{k})=\log(\beta_{k}/\alpha_{k})$. By the Birkhoff-Bushell theorem,
@@
Proposition 3, then, follows immediately from Lemma 5 and Proposition 6. ∎
A.5. Proof for Proposition 4
@@
$f_{E_{2}}(\mathcal{Q}_{1})\subseteq f_{E_{2}}(\mathcal{Q}_{2})\subseteq f_{E_{2}}(\mathcal{Q}_{3}).$ (4)
@@
$\underline{f}(\mathcal{Q}_{1})\subseteq f_{E_{2}}(\mathcal{Q}_{1})\quad\text{ and }\quad f_{E_{2}}(\mathcal{Q}_{3})\subseteq\overline{f}(\mathcal{Q}_{3}).$ (5)
@@
Now, assuming that the conditions of Corollary 2.2 hold for both $\underline{f}$ and $\overline{f}$, and that the Hausdorff limit $\lim_{n\rightarrow\infty}\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})$ exists, we can conclude that
@@
thus proving the statement. ∎
Appendix B Additional Result: Finitely Additive Case

### v2

@@ A.5. Classical Fixed Point Theorems
The reader may ask themselves why, in Theorem 3, we chose to prove a generalization of Boyd-Wong’s fixed point theorem; In other words, one might ask whether the convergence of the orbit $(\mathcal{P}_{n})$ could be established under weaker assumptions.
First, let us discuss $\mathcal{X}$ being Polish. Notice that if $\mathcal{X}$ is not separable, then $\Delta_{\mathcal{X}}$ may not be metrizable, which in turn means that we cannot talk about the Hausdorff metric $d_{H}$ on $\mathscr{C}$, let alone about Hausdorff convergence. Also, if $\mathcal{X}$ is not complete, then there might be an orbit $(\mathcal{P}_{n})$ of some $\mathcal{P}_{0}\in\mathscr{C}$ that has an accumulation point, but does not converge to it.
@@
Appendix B Proofs and Additional Results
B.1. Proof for Theorem 1
@@
Brouwer-Schauder-Tychonoff fixed point theorem (Aliprantis and Border,, 2006, Theorem 17.56) guarantees a point $h_{*}\in K$ with $\tilde{f}(h_{*})=h_{*}$, since $K$ is compact and convex in the locally convex topological (Hausdorff) vector space $Z$.
@@
Let $P_{*}\coloneqq J^{-1}(h_{*})$. Then $f(P_{*})=J^{-1}\bigl(\tilde{f}(J(P_{*}))\bigr)=P_{*}$, so $P_{*}$ is a fixed point of $f$. Finally, the collection of fixed points of $f$, $\mathrm{Fix}(f)=\{P\in\mathscr{C}:d_{H}(P,f(P))=0\}$, is closed in compact $\mathscr{C}$, hence compact. $\blacksquare$
B.2. Proof for Proposition 2
@@
Notice now that, at the beginning of the proof, we picked a generic prior $P^{\text{ex}}\in\text{ex}\mathcal{P}$ and a generic likelihood $L_{k}\in\text{ex}\mathcal{L}$. This means that the functions $A\mapsto g_{k}(P^{\text{ex}})(A)$ are continuous, for all $k$ and all $P^{\text{ex}}$. Then, since $\Delta_{\Theta}$ is locally convex (as a result of $\Theta$ being compact metric), we conclude that $f$ is Hausdorff continuous by Aliprantis and Border, (2006, Theorem 17.37). $\blacksquare$
B.3. Proof of Corollary 2.1
From the proof of Theorem 1, we know that, under our assumptions, $(\mathscr{C},d_{H})$ is a Hausdorff compact metric space. The result, then, follows from Edelstein’s fixed point theorem, Theorem 7 in Appendix A.5. $\blacksquare$
B.4. Proof for Theorem 3
@@
Thus, $C\in\mathscr{C}$, proving that $\mathscr{C}$ is a closed subset of $\mathscr{F}$. Since $\mathscr{F}$ is complete under $d_{H}$, and $\mathscr{C}$ is a closed subset of $\mathscr{F}$, it follows that $\mathscr{C}$ is complete under $d_{H}$.
Since $(\mathscr{C},d_{H})$ is a compact metric space, then it is also complete. Our claim, then, follows from Boyd-Wong’s fixed point theorem, Theorem 8 in Appendix A.5. $\blacksquare$
Remark 4 (Information about Convergence Rate).
Notice that the only information we get from Theorem 3 about the convergence rate of $(\mathcal{P}_{n})$ is that it is sublinear and governed by $\psi$. For example, if $\psi$ satisfies additional properties—e.g. Lipschitzianity or $\psi(t)=qt$, with $q\in(0,1)$—the rate becomes tractable, as the setting becomes similar to Banach’s. It is also worth mentioning that there is an active literature studying the convergence speed for Boyd-Wong-type contractions under extra assumptions on the structure of $\psi$ (Gupta and Mansotra,, 2023; Gautam and Kaur,, 2023; Singh et al.,, 2022; Das et al.,, 2024).
B.5. Proof for Proposition 4
@@
Multiplication by $\ell_{k}(E\mid\cdot)$ defines a positive linear map $M_{k}:p\mapsto\ell_{k}(E\mid\cdot)p$ on the positive cone of $L^{1}(\mu)$; the normalization $p\mapsto p/\int p\text{d}\mu$ is projective (leaves $d_{\mathrm{H}}$ unchanged). Since $\alpha_{k}\leq\ell_{k}\leq\beta_{k}$ a.e., the projective diameter of $M_{k}$ equals $\mathscr{D}(M_{k})=\log(\beta_{k}/\alpha_{k})$. By the Birkhoff-Bushell theorem,
@@
Proposition 4, then, follows immediately from Lemma 9 and Proposition 10. $\blacksquare$
B.6. Proof for Proposition 5
@@
(5) $f_{E_{2}}(\mathcal{Q}_{1})\subseteq f_{E_{2}}(\mathcal{Q}_{2})\subseteq f_{E_{2}}(\mathcal{Q}_{3}).$
@@
(6) $\underline{f}(\mathcal{Q}_{1})\subseteq f_{E_{2}}(\mathcal{Q}_{1})\quad\text{ and }\quad f_{E_{2}}(\mathcal{Q}_{3})\subseteq\overline{f}(\mathcal{Q}_{3}).$
@@
Now, assuming that the conditions of Corollary 3 hold for both $\underline{f}$ and $\overline{f}$, and that the Hausdorff limit $\lim_{n\rightarrow\infty}\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})$ exists, we can conclude that
@@
thus proving the statement. $\blacksquare$
B.7. Two sufficient conditions for existence of the Hausdorff limit
In this section, we show that the two conditions (a) and (b) in Remark 2 are indeed sufficient for the Hausdorff limit $\lim_{n\rightarrow\infty}\bigcirc_{i=1}^{n}f_{E_{i}}(\mathcal{P}_{0})$ to exist. We prove a general version of the results; it is then enough to put $X=\Delta_{\mathcal{X}}$ and $\mathscr{C}$ as our class of credal sets to retrieve our setting.
Let $(X,d)$ be a compact metric space and let $\mathcal{K}(X)$ denote the set of all nonempty compact subsets of $X$, endowed with the Hausdorff metric $d_{H}$. Let $\mathscr{C}\subseteq\mathcal{K}(X)$ and let $(f_{i})_{i\geq 1}$ be a sequence of maps $f_{i}:\mathscr{C}\to\mathscr{C}$. Fix $\mathcal{P}_{0}\in\mathscr{C}$ and define
$\mathcal{P}_{n}\coloneqq f_{n}\circ f_{n-1}\circ\cdots\circ f_{1}(\mathcal{P}_{0}),\qquad n\geq 1.$
Then the following are sufficient for the existence of the Hausdorff limit $\lim_{n\to\infty}\mathcal{P}_{n}$ in $(\mathcal{K}(X),d_{H})$:
  • (a)
    Uniform Hausdorff contraction. There exists a constant $L\in[0,1)$ such that, for all $i\geq 1$ and all $\mathcal{P},\mathcal{Q}\in\mathscr{C}$,
    $d_{H}\bigl(f_{i}(\mathcal{P}),f_{i}(\mathcal{Q})\bigr)\leq Ld_{H}(\mathcal{P},\mathcal{Q}).$
  • (b)
    Nested chain with vanishing diameters. The sequence $(\mathcal{P}_{n})_{n\geq 1}$ is nested and decreasing, i.e. $\mathcal{P}_{n+1}\subseteq\mathcal{P}_{n}$ for all $n\geq 1$, and its diameters satisfy
    $\operatorname{diam}(\mathcal{P}_{n})\coloneqq\sup\{d(x,y):x,y\in\mathcal{P}_{n}\}\longrightarrow 0\quad\text{as }n\to\infty.$
In case (a), the sequence $(\mathcal{P}_{n})$ is Cauchy in $(\mathcal{K}(X),d_{H})$ and hence convergent. In case (b), the sequence $(\mathcal{P}_{n})$ converges in $d_{H}$ to a singleton $\{x_{\star}\}\in\mathcal{K}(X)$.
Proof.
It is well known that if $(X,d)$ is compact, then $(\mathcal{K}(X),d_{H})$ is a compact metric space (hence complete). We use this fact in both parts.
Proof of (a). Assume that there exists $L\in[0,1)$ such that
$d_{H}\bigl(f_{i}(\mathcal{P}),f_{i}(\mathcal{Q})\bigr)\leq Ld_{H}(\mathcal{P},\mathcal{Q})\quad\text{for all }i\geq 1,\mathcal{P},\mathcal{Q}\in\mathscr{C}.$
Define $\mathcal{P}_{n}$ as in the statement and let
$\delta_{n}\coloneqq d_{H}(\mathcal{P}_{n},\mathcal{P}_{n-1}),\qquad n\geq 1.$
Then, for $n\geq 1$,
$\delta_{n+1}=d_{H}\bigl(\mathcal{P}_{n+1},\mathcal{P}_{n}\bigr)=d_{H}\bigl(f_{n+1}(\mathcal{P}_{n}),f_{n+1}(\mathcal{P}_{n-1})\bigr)\leq Ld_{H}(\mathcal{P}_{n},\mathcal{P}_{n-1})=L\,\delta_{n}.$
By induction this yields
$\delta_{n+1}\leq L^{n}\delta_{1}=L^{n}d_{H}(\mathcal{P}_{1},\mathcal{P}_{0}),\qquad n\geq 0.$
Let $m>n\geq 0$. Using the triangle inequality and the above bound,
$d_{H}(\mathcal{P}_{m},\mathcal{P}_{n})\leq\sum_{j=n}^{m-1}d_{H}(\mathcal{P}_{j+1},\mathcal{P}_{j})=\sum_{j=n}^{m-1}\delta_{j+1}\leq\sum_{j=n}^{m-1}L^{j}\delta_{1}=L^{n}\frac{1-L^{m-n}}{1-L}\delta_{1}\leq\frac{L^{n}}{1-L}\delta_{1}.$
Since $L\in[0,1)$, the right-hand side tends to 0 as $n\to\infty$, uniformly in $m>n$. Thus $(\mathcal{P}_{n})$ is a Cauchy sequence in the complete metric space $(\mathcal{K}(X),d_{H})$, and hence there exists $\mathcal{P}_{\star}\in\mathcal{K}(X)$ such that $d_{H}(\mathcal{P}_{n},\mathcal{P}_{\star})\to 0$ as $n\to\infty$.
Proof of (b). Assume now that $(\mathcal{P}_{n})$ is nested and decreasing (i.e. $\mathcal{P}_{n+1}\subseteq\mathcal{P}_{n}$ for all $n$) and that $\operatorname{diam}(\mathcal{P}_{n})\to 0$ as $n\to\infty$.
First, we show that $\bigcap_{n=1}^{\infty}\mathcal{P}_{n}$ is a singleton. Pick any sequence $(x_{n})_{n\geq 1}$ such that $x_{n}\in\mathcal{P}_{n}$ for each $n$. For $m>n$ we have $x_{n},x_{m}\in\mathcal{P}_{n}$, because $\mathcal{P}_{m}\subseteq\mathcal{P}_{n}$. Hence
$d(x_{n},x_{m})\leq\operatorname{diam}(\mathcal{P}_{n}).$
As $n\to\infty$ we have $\operatorname{diam}(\mathcal{P}_{n})\to 0$, so $(x_{n})$ is a Cauchy sequence in the complete space $X$, and thus converges to some $x_{\star}\in X$.
Next, for each fixed $n$, the tail $\{x_{m}:m\geq n\}$ is contained in $\mathcal{P}_{n}$, which is closed. Since $x_{m}\to x_{\star}$, it follows that $x_{\star}\in\mathcal{P}_{n}$ for every $n$, so that $x_{\star}\in\bigcap_{n=1}^{\infty}\mathcal{P}_{n}$ and the intersection is nonempty.
To see uniqueness, let $y_{\star}\in\bigcap_{n=1}^{\infty}\mathcal{P}_{n}$ be arbitrary. Then $x_{\star},y_{\star}\in\mathcal{P}_{n}$ for every $n$, hence
$d(x_{\star},y_{\star})\leq\operatorname{diam}(\mathcal{P}_{n})\quad\text{for all }n.$
Letting $n\to\infty$ we obtain $d(x_{\star},y_{\star})=0$, so $x_{\star}=y_{\star}$. Thus $\bigcap_{n=1}^{\infty}\mathcal{P}_{n}=\{x_{\star}\}$ is a singleton.
Finally, we show that $\mathcal{P}_{n}\to\{x_{\star}\}$ in the Hausdorff metric. For each $n$,
$d_{H}(\mathcal{P}_{n},\{x_{\star}\})=\max\left\{\sup_{x\in\mathcal{P}_{n}}d(x,x_{\star}),\sup_{y\in\{x_{\star}\}}\inf_{x\in\mathcal{P}_{n}}d(y,x)\right\}.$
Since $x_{\star}\in\mathcal{P}_{n}$ for all $n$, the second term is zero, and we have
$d_{H}(\mathcal{P}_{n},\{x_{\star}\})=\sup_{x\in\mathcal{P}_{n}}d(x,x_{\star})\leq\operatorname{diam}(\mathcal{P}_{n}).$
As $\operatorname{diam}(\mathcal{P}_{n})\to 0$, it follows that $d_{H}(\mathcal{P}_{n},\{x_{\star}\})\to 0$ as $n\to\infty$. Hence $\mathcal{P}_{n}$ converges in the Hausdorff metric to the singleton $\{x_{\star}\}$. ∎
Appendix C Simulation Details
To reproduce Figure 2, one could run the following script:
Listing 1: Iterative Bayesian Updating Simulation
1# Initialisation
2N = 50
3NUM_ROUNDS = 150
4
5# Data Generating Process
6TRUE_SIGMA, TRUE_MEAN = 2.0, 0.0
7theta = np.random.normal(TRUE_MEAN, TRUE_SIGMA)
8samples = np.random.normal(theta, scale=1.0, size=(N))
9
10# Set up prior parameter sets
11PRIOR_MEANS = [2.0, 3.0, -2.0]
12PRIOR_SIGMAS = [1.0, 2.0, 2.0]
13
14posterior_means_set, posterior_SIGMAs_set = [], []
15for i in range(len(PRIOR_MEANS)):
16 PRIOR_SIGMA = PRIOR_SIGMAS[i]
17 PRIOR_MEAN = PRIOR_MEANS[i]
18
19 posterior_means, posterior_SIGMAs = [], []
20 for t in range(NUM_ROUNDS):
21 posterior_sigma = 1/(1/PRIOR_SIGMA + N/TRUE_SIGMA)
22 posterior_mean = posterior_sigma * (
23 PRIOR_MEAN/PRIOR_SIGMA + N * samples.mean()/TRUE_SIGMA
24 )
25
26 PRIOR_MEAN = posterior_mean
27 PRIOR_SIGMA = posterior_sigma
28
29 posterior_means.append(posterior_mean)
30 posterior_SIGMAs.append(posterior_sigma)
31
32 posterior_means_set.append(np.array(posterior_means))
33 posterior_SIGMAs_set.append(np.array(posterior_SIGMAs))
34
35consecutive_diff_in_mean = 0
36consecutive_diff_in_SIGMA = 0
37for j in range(3):
38 consecutive_diff_in_mean += np.abs(
39 posterior_means_set[j][:-1] - posterior_means_set[0][1:]
40 )
41 consecutive_diff_in_SIGMA += np.abs(
42 posterior_SIGMAs_set[j][:-1] - posterior_SIGMAs_set[0][1:]
43 )
44
45consecutive_diff_in_mean /= 3
46consecutive_diff_in_SIGMA /= 3
47
48plt.figure(figsize=(4.5, 3))
49plt.plot(range(NUM_ROUNDS-1), consecutive_diff_in_SIGMA,
50 label="Consecutive parameter difference", linewidth=3)
51plt.xlabel("Number of Rounds \$t\$")
[…]
53plt.legend()
54plt.show()
Appendix D Additional Result: Finitely Additive Case
The reason why the results in Section 3 may not hold in the finitely additive case is that, if $\mathcal{X}$ is not finite, then the space of finitely additive probabilities may not be metrizable, and so we cannot talk about Hausdorff distance, which played a crucial role in the proofs in Appendix B. On the other hand, the space of countably additive probability measures (under the weak topology) is always metrizable, provided $\mathcal{X}$ is metrizable. These concerns should not worry the practitioner, though, as classical probabilistic ML works with countably additive probability measures.